# Theoretical modeling of approximate universality of tidally deformed neutron stars

### v1

@@ I.1 Neutron star physics
NSs are the densest objects in the Universe, offering a unique opportunity to explore the properties of matter at supernuclear densities Lattimer and Prakash (2016); Özel and Freire (2016); Burgio et al. (2021). The central density is a few times the nuclear saturation density $\rho_{\rm nuc}=2.8\times 10^{14}~{}{\rm g~{}cm^{-3}}$, a regime that cannot yet be accessed from the first principles. Determining the equation of state (EoS) of nuclear matter in an extreme environment is one of the outstanding goals in nuclear physics. The NS mass-radius relation is in one-to-one correspondence with the EoS. Therefore, independent accurate measurements of the NS mass and radius with multifaceted observations allow us to determine the EoS. NS masses have been accurately measured through the Shapiro time delay (and other post-Keplerian parameters) in the arrival times of radio pulses emitted by millisecond pulsars Jacoby et al. (2005); Verbiest et al. (2008); Demorest et al. (2010). NS radii, together with masses, are determined by the NICER mission using X-ray that is emitted from hot spots on the surfaces of millisecond pulsars Riley et al. (2019); Miller et al. (2019, 2021); Choudhury et al. (2024). GW observations of binary NSs allow us to constrain the EoS and radii by measuring the stellar tidal deformability Abbott et al. (2017a, 2018, 2019, 2020).
@@ I.3 Summary of this work
We conduct a comprehensive, theoretical study on the approximate universality of NSs. Our approach employs a semi-analytic relativistic stellar interior model, called a modified Tolman VII model Posada et al. (2022); Jiang and Yagi (2019), and solve the tidal perturbations based on the post-Minkowskian expansion, thereby deriving the universal Love relations among the tidal deformabilities of nonrotating NSs:
@@
The suppression mechanism is continuously reinforced as the local adiabatic index increases. Our findings suggest that the approximate universality of perturbed NSs can be attributed to low compressibility. Indeed, we find that stellar configurations in equilibrium tend to become insensitive to variations in EoSs as the local adiabatic index increases. These observations suggest the following picture: incompressible stars correspond to a restricted subset of equilibrium stellar configurations in the phase space spanned by physical degrees of freedom, as the energy density remains constant throughout the entire interior. Linear perturbations to such incompressibe stars induce the strongest possible responses, setting upper bounds on their magnitudes. Stars deviate from this limit by decreasing the adiabatic index. For NSs, stellar configurations remain insensitive to variations in the EoSs due to low compressibility. A decrease in the adiabatic index reduces the magnitudes of the responses in a manner shared among the perturbations, leading to approximate universality. Eventually, as the adiabatic index decreases further, the degrees of freedom associated with an EoS expand, and in turn, the similarity of the responses is lost, resulting in the loss of universality.
@@ III.1 Tidal deformability
The tidal perturbation is decomposed into the electric-type and magnetic-type sectors, subject to the parity transformation Hinderer (2008); Damour and Nagar (2009); Binnington and Poisson (2009). With harmonic decomposition, the tidal deformation of the star is quantified by a set of the dimensionless $\ell$th electric-type and magnetic-type tidal deformabilities $\bar{\lambda}_{\ell}$ and $\bar{\sigma}_{\ell}$ Damour and Nagar (2009); Yagi (2014)
@@ III.2.1 Even–Odd Love relations
Figure  4: (Top panel) The $\bar{\lambda}_{2}-\bar{\sigma}_{2}$ relation. The black line is the power-law relation (1). The magenta dashed line is the empirical fit for realistic EoSs in Yagi (2014). The colored circles denote a set of $(\bar{\lambda}_{2},\bar{\sigma}_{2})$ numerically computed for various $\alpha$ in the modified Tolman VII model given discrete compactness. Observe that the variation of the results with respect to $\alpha$ is indistinguishable. (Bottom panel)  The fractional difference between the numerical values represented by the circles and the theoretical power-law relation (1) (solid curves). The black dashed curve denotes the fractional difference between the theoretical fitting expressions. Note that the theoretical relation is approximately derived for the Tolman VII solution ($\alpha=1$), leading to the smallest fractional difference in a wide range of $\bar{\lambda}_{2}$.
Figure 4 presents the $\bar{\lambda}_{2}-\bar{\sigma}_{2}$ relation, demonstrating the insensitivity of the results with respect to variation in $\alpha$ and the excellent agreement of the power-law expression (1) with the previous empirical finding by Ref. Yagi (2014). Note that the colored circles denote a set of $(\bar{\lambda}_{2},\bar{\sigma}_{2})$ numerically computed for various $\alpha$ on the background described by the modified Tolman VII solution.
@@ IV.1 EoS independence of the power-law exponent
Equations (13) and (14) imply that $\bar{\lambda}_{\ell}$ and $\bar{\sigma}_{\ell}$ scale as ${\cal C}^{-2\ell-1}$ and ${\cal C}^{-2\ell}$, respectively, to leading order of the PM expansion. A comparison by eliminating ${\cal C}$ from them reveals the scaling relation, $\bar{\sigma}_{\ell}\propto\bar{\lambda}_{\ell}^{2\ell/(2\ell+1)}$, which determines the exponent in Eq. (16) for $\ell=2$. Similarly, one finds $\bar{\lambda}_{\ell^{\prime}}\propto\bar{\lambda}_{\ell}^{(2\ell^{\prime}+1)/(2\ell+1)}$ for $\ell^{\prime}(\neq\ell)$, explaining the exponent in Eq. (19) for $(\ell,\ell^{\prime})=(2,3)$. The exponent is thus determined by the scaling law of the tidal deformablities with respect to ${\cal C}$, which is independent of the underlying microphysics.
@@ IV.4 Universality from low compressibility
We find that the trend in the $\alpha$ dependence of the power-law coefficients is inherited from that of the equilibrium stellar configuration, i.e., the modified Tolman VII solution. Figure 8 presents the fractional difference in the energy density profiles (8) for similar values of $\alpha$ with a fixed $\bar{\lambda}_{2}=1000$, showing that the difference is typically less than $10\%$ in a large potion of the interior and increases for larger $\alpha$. Similar behavior can be found in other background quantities, such as $p$, $\nu$, and $m$. In other words, the stellar configuration in equilibrium tends to become insensitive to $\alpha$ variations as $\alpha$ approaches zero.
@@ V Discussion and extensions
From a practical perspective, if the universal Love relations hold even for rotational TLNs, they could provide a powerful tool for waveform modeling. This is because the leading-order rotational TLN enters at $6.5$PN order, which is lower than the $\ell\geq 3$th electric-type and magnetic-type tidal deformabilities Jiménez Forteza et al. (2018). Furthermore, neglecting the spin-tidal coupling may introduce a significant bias in parameter estimation for NSs with dimensionless spin of $\chi\sim 0.1$ with third-generation detectors Jiménez Forteza et al. (2018); Castro et al. (2022). Establishing an EoS insensitive connection between the rotational TLN and the quadrupolar electric-type tidal deformability could enhance the ability to infer the EoS at supernuclear densities and test GR in the strong-field regime.
@@ Acknowledgements.
T.K. and G.R.M. acknowledge support from the Villum Investigator program supported by the VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. This work has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101131233. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. 184. K.Y. acknowledges support from NSF Grant No. PHY-2309066, No. PHYS-2339969, and the Owens Family Foundation.
@@ A.2 Fully analytic model
ϕ=italic-$\phi=$ $C_{2}-\frac{1}{2}\ln\left(\xi^{2}-\frac{5}{6}+\sqrt{\frac{5e^{-\lambda}}{8\pi\rho_{c}R^{2}}}\right),$
@@ B.1.1 Stellar perturbations
The even-parity metric perturbation induces the displacement of $\rho$, $p$, and the four velocity. We parameterize the Lagrangian displacement of the fluid variables by two functions, $W=W(r)$ and $V=V(r)$, so that the components of the displacement vector $\xi^{\mu}$ are expressed by Lindblom and Detweiler (1983)
@@
The Eulerian variation of the density and pressure at linear order is related to the change of the baryon number density as follows. First, we assume that the baryon number of each fluid element is conserved to linear order:
$\nabla_{\mu}\left[\left(n+\Delta n\right)\left(u^{\mu}+\Delta u^{\mu}\right)\right]=0,$ (38)
@@
The EoS links the pressure and energy density at the background level. In general, the Lagrangian perturbation to pressure, $\Delta p$, depends on the EoS in equilibrium configurations and on chemical reactions during the dynamical process through $\gamma$, which correlates with the entropy per baryon and the change in the fractional abundance of nuclear species inside the NS. If the relaxation timescale of perturbed configurations toward equilibrium is much shorter than the orbital period, the reactions may maintain equilibrium even in perturbed configurations in the adiabatic manner. The fractional abundance of nuclear species then shifts continuously while maintaining equilibrium, and hence, the variation in pressure with respect to that in density in perturbations approximately follows the same relation as in equilibrium configurations. In such a situation, the EoS for the background also holds for perturbed configurations. The connection between density and pressure perturbations is provided by $\gamma=\gamma(\rho)$ through
@@ B.1.2 Perturbation equations
Now, let us linearize Einstein’s equations and the matter conservation law, $\delta G_{\mu\nu}=8\pi\delta T_{\mu\nu}$ and $\nabla_{\mu}\delta T^{\mu\nu}=0$. Combining various components in these equations, we derive a set of perturbation equations for $H,W$, and $V$ as follows.
@@
Next, we look at the perturbed stress-energy conservation equation. From $\nabla_{\mu}T^{\mu}_{~{}\theta}=0$, one derives the expression for $W^{\prime}$ in terms of $(H,W,H^{\prime},V^{\prime})$, leading to $\nabla_{\mu}T^{\mu}_{~{}\varphi}=0$. Next, $\nabla_{\mu}T^{\mu}_{~{}r}=0$ gives $V^{\prime}$ in terms of $(H,W,V,H^{\prime})$, thereby allowing one to eliminate $V^{\prime}$ from $W^{\prime}$ and $H^{\prime\prime}$.
@@
$W^{\prime}=\alpha_{W,H^{\prime}}H^{\prime}+\alpha_{W,H}H+\alpha_{W,W}W++\alpha_{W,V}V,$ (45)
@@ B.1.6 Electric-type tidal deformability
$\left.\frac{\mathbb{I}_{H}}{\mathbb{E}_{H}}=-\frac{H_{T}^{\prime}-y_{\rm in}^{+}H_{T}}{H_{R}^{\prime}-y_{\rm in}^{+}H_{R}}\right|_{r=R}.$ (57)
@@
$\bar{\lambda}_{\ell}=\frac{2}{\left(2\ell-1\right)!!}{\cal C}^{-2\ell-1}\frac{1}{{\cal C}^{2\ell+1}}\frac{p_{\ell,0}^{+}+p_{\ell,1}^{+}{\cal C}+p_{\ell,2}^{+}{\cal C}^{2}+p_{\ell,3}^{+}{\cal C}^{3}}{q_{\ell,0}^{+}+q_{\ell,1}^{+}{\cal C}+q_{\ell,2}^{+}{\cal C}^{2}+q_{\ell,3}^{+}{\cal C}^{3}},$ (58)
@@ B.2.1 Stellar perturbation equations
$h_{\mu\nu}^{-}dx^{\mu}dx^{\nu}=2h^{\ell m}S_{\theta}^{\ell m}dtd\theta+2h^{\ell m}S_{\varphi}^{\ell m}dtd\varphi$
$+2h_{1}^{\ell m}S_{\theta}^{\ell m}drd\theta+2h_{1}^{\ell m}S_{\varphi}^{\ell m}drd\varphi.$ (60)
@@
In the odd-parity sector, the metric perturbation is not coupled to the pressure and density perturbations but induces the displacement of the four velocity. The Eulerian perturbation of the four velocity is parameterized by a function $U=U(r)$ Price and Thorne (1969); Pani et al. (2015b, 2018)
$\delta u^{t}=$ $0,\quad\delta u^{r}=0,\quad\delta u^{A}=\frac{Ue^{\nu/2}}{4\pi r^{2}\left(\rho+p\right)}S^{A}e^{-i\omega t}.$ (61)
@@
One finds from $\nabla_{\mu}\delta T^{\mu}_{~{}\theta}=0$ that

### v2

@@ I.1 Neutron star physics
NSs are the densest objects in the Universe, offering a unique opportunity to explore the properties of matter at supranuclear densities Lattimer and Prakash (2016); Özel and Freire (2016); Burgio et al. (2021). The central density is a few times the nuclear saturation density $\rho_{\rm nuc}=2.8\times 10^{14}~{}{\rm g~{}cm^{-3}}$, a regime that cannot yet be accessed from the first principles. Determining the equation of state (EoS) of nuclear matter in an extreme environment is one of the outstanding goals in nuclear physics. The NS mass-radius relation is in one-to-one correspondence with the EoS. Therefore, independent accurate measurements of the NS mass and radius with multifaceted observations allow us to determine the EoS. NS masses have been accurately measured through the Shapiro time delay (and other post-Keplerian parameters) in the arrival times of radio pulses emitted by millisecond pulsars Jacoby et al. (2005); Verbiest et al. (2008); Demorest et al. (2010). NS radii, together with masses, are determined by the NICER mission using X-ray that is emitted from hot spots on the surfaces of millisecond pulsars Riley et al. (2019); Miller et al. (2019, 2021); Choudhury et al. (2024). GW observations of binary NSs allow us to constrain the EoS and radii by measuring the stellar tidal deformability Abbott et al. (2017a, 2018, 2019, 2020).
@@ I.3 Summary of this work
We conduct a comprehensive, theoretical study on the approximate universality of NSs. Our approach employs a semi-analytic relativistic stellar interior model, referred to as a modified Tolman VII model Posada et al. (2022); Jiang and Yagi (2019), and solves the tidal perturbations based on the post-Minkowskian expansion. We semi-analytically derive the universal Love relations among the tidal deformabilities of nonrotating NSs by solving tidal perturbations on the Tolman VII solution Tolman (1939), which is a particular case of the modified Tolman VII model:
@@
The suppression mechanism is continuously reinforced as the local adiabatic index increases. Our findings suggest that the approximate universality of perturbed NSs can be attributed to low compressibility. Indeed, we find that stellar configurations in equilibrium tend to become insensitive to variations in EoSs as the local adiabatic index increases. These observations suggest the following picture: incompressible stars correspond to a restricted subset of equilibrium stellar configurations in the phase space spanned by physical degrees of freedom, as the energy density remains constant throughout the entire interior. Linear perturbations to such incompressible stars induce the strongest possible responses, setting upper bounds on their magnitudes. Stars deviate from this limit by decreasing the adiabatic index. For NSs, stellar configurations remain insensitive to variations in the EoSs due to low compressibility. A decrease in the adiabatic index reduces the magnitudes of the responses in a manner shared among the perturbations, leading to approximate universality. Eventually, as the adiabatic index decreases further, the degrees of freedom associated with an EoS expand, and in turn, the similarity of the responses is lost, resulting in the loss of universality.
@@ III.1 Tidal deformability
The tidal perturbation is decomposed into the electric-type and magnetic-type sectors, subject to the parity transformation Hinderer (2008); Damour and Nagar (2009); Binnington and Poisson (2009). With harmonic decomposition, the tidal deformation of the star is quantified by a set of the dimensionless $\ell$th electric-type and magnetic-type tidal deformabilities $\bar{\lambda}_{\ell}$ and $\bar{\sigma}_{\ell}$ Damour and Nagar (2009); Yagi (2014),
@@ III.2.1 Even–Odd Love relations
Figure  4: (Top panel) The $\bar{\lambda}_{2}-\bar{\sigma}_{2}$ relation. The black line is the power-law relation (1). The magenta dashed line is the empirical fit for realistic EoSs in Yagi (2014). The colored circles denote a set of $(\bar{\lambda}_{2},\bar{\sigma}_{2})$ numerically computed for various $\alpha$ in the modified Tolman VII model without relying on the PM expansion, given discrete compactness. Observe that the variation of the results with respect to $\alpha$ is indistinguishable. (Bottom panel)  The fractional difference between the numerical values represented by the circles and the theoretical power-law relation (1) (solid curves). The black dashed curve denotes the fractional difference between the theoretical and fitting expressions. Note that the theoretical relation is approximately derived for the Tolman VII solution ($\alpha=1$) within the PM expansion, leading to the smallest fractional difference in a wide range of $\bar{\lambda}_{2}$.
Figure 4 presents the $\bar{\lambda}_{2}-\bar{\sigma}_{2}$ relation, demonstrating the insensitivity of the results with respect to variation in $\alpha$ and the excellent agreement of the power-law expression (1) with the previous empirical finding by Ref. Yagi (2014). [^6] Note that the bottom panel does not display the relative error between the semi-analytically constructed curves across various $\alpha$ and the theoretical relation based on $\alpha=1$ (see the caption in Fig. 4). The finite difference for $\alpha=1$ merely reflects the accuracy of the PM expansion in that case. The same remark applies to Figs. 51213, and 14. Note that the colored circles denote a set of $(\bar{\lambda}_{2},\bar{\sigma}_{2})$ numerically computed for various $\alpha$ on the background described by the modified Tolman VII solution.
@@ IV.1 EoS independence of the power-law exponent
Equations (13) and (14) imply that $\bar{\lambda}_{\ell}$ and $\bar{\sigma}_{\ell}$ scale as ${\cal C}^{-2\ell-1}$ and ${\cal C}^{-2\ell}$, respectively, to leading order of the PM expansion. A comparison by eliminating ${\cal C}$ from them reveals the scaling relation, $\bar{\sigma}_{\ell}\propto\bar{\lambda}_{\ell}^{2\ell/(2\ell+1)}$, which determines the exponent in Eq. (16) for $\ell=2$. Similarly, one finds $\bar{\lambda}_{\ell^{\prime}}\propto\bar{\lambda}_{\ell}^{(2\ell^{\prime}+1)/(2\ell+1)}$ for $\ell^{\prime}(\neq\ell)$, explaining the exponent in Eq. (19) for $(\ell,\ell^{\prime})=(2,3)$. The exponent is thus determined by the scaling law of the tidal deformabilities with respect to ${\cal C}$, which is independent of the underlying microphysics.
@@ IV.4 Universality from low compressibility
We find that the trend in the $\alpha$ dependence of the power-law coefficients is inherited from that of the equilibrium stellar configuration, i.e., the modified Tolman VII solution. Figure 8 presents the fractional difference in the energy density profiles (8) for similar values of $\alpha$ with a fixed $\bar{\lambda}_{2}=1000$, showing that the difference is typically less than $10\%$ in a large potion of the interior and increases for larger $\alpha$. Similar behavior can be found in other background quantities, such as $p$, $\nu$, and $m$. In other words, the stellar configuration in equilibrium tends to become insensitive to $\alpha$ variations as $\alpha$ approaches zero. The “valley” observed in Fig. 8 originates from the crossing of the energy-density profiles for different values of $\alpha$ around $\xi=0.6$, as shown in Fig. 2.
@@ V Discussion and extensions
From a practical perspective, if the universal Love relations hold even for rotational TLNs, they could provide a powerful tool for waveform modeling. This is because the leading-order rotational TLN enters at $6.5$PN order, which is lower than the $\ell\geq 3$rd electric-type and magnetic-type tidal deformabilities Jiménez Forteza et al. (2018). Furthermore, neglecting the spin-tidal coupling may introduce a significant bias in parameter estimation for NSs with dimensionless spin of $\chi\sim 0.1$ with third-generation detectors Jiménez Forteza et al. (2018); Castro et al. (2022). Establishing an EoS insensitive connection between the rotational TLN and the quadrupolar electric-type tidal deformability could enhance the ability to infer the EoS at supernuclear densities and test GR in the strong-field regime.
@@ Acknowledgements.
We are indebted to Zexin Hu and Lijing Shao for their valuable comments and insights. T.K. and G.R.M. acknowledge support from the Villum Investigator program supported by the VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. This work has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101131233. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. 184. K.Y. acknowledges support from NSF Grant No. PHY-2309066, No. PHYS-2339969, and the Owens Family Foundation.
@@ A.2 Fully analytic model
ϕ=italic-$\phi=$ $C_{2}-\frac{1}{2}\ln\left(\xi^{2}-\frac{5}{6}+\sqrt{\frac{5e^{-\lambda|_{\alpha=1}}}{8\pi\rho_{c}R^{2}}}\right),$
@@ B.1.1 Stellar perturbations
The even-parity metric perturbation induces the displacement of $\rho$, $p$, and the four velocity. We parametrize the Lagrangian displacement of the fluid variables by two functions, $W=W(r)$ and $V=V(r)$, so that the components of the displacement vector $\xi^{\mu}$ are expressed by Lindblom and Detweiler (1983)
@@
The Eulerian variation of the density and pressure at linear order is related to the change of the baryon number density as follows. First, we assume that the baryon number of each fluid element in the perturbed configuration is conserved:
$\Delta\left[\nabla_{\mu}\left(nu^{\mu}\right)\right]=0,$ (38)
@@
The EoS links the pressure and energy density at the background level. In general, the Lagrangian perturbation to pressure, $\Delta p$, depends on the EoS in equilibrium configurations and on chemical reactions during the dynamical process, which correlates with the entropy per baryon and the change in the fractional abundance of nuclear species inside the NS. If the relaxation timescale of perturbed configurations toward equilibrium is much shorter than the orbital period, the reactions may maintain equilibrium even in perturbed configurations in the adiabatic manner. The fractional abundance of nuclear species then shifts continuously while maintaining equilibrium, and hence, the variation in pressure with respect to that in density in perturbations approximately follows the same relation as in equilibrium configurations. In such a situation, the EoS for the background also holds for perturbed configurations. The connection between density and pressure perturbations is provided by $\gamma=\gamma(\rho)$ through
@@ B.1.2 Perturbation equations
Now, let us linearize Einstein’s equations and the matter conservation law, $\delta G_{\mu\nu}=8\pi\delta T_{\mu\nu}$ and $\delta(\nabla_{\mu}T^{\mu\nu})=0$. Combining various components in these equations, we derive a set of perturbation equations for $H,W$, and $V$ as follows.
@@
Next, we look at the perturbed stress-energy conservation equation. From $\delta(\nabla_{\mu}T^{\mu}_{~{}\theta})=0$, one derives the expression for $W^{\prime}$ in terms of $(H,W,H^{\prime},V^{\prime})$, leading to $\delta(\nabla_{\mu}T^{\mu}_{~{}\varphi})=0$. Next, $\delta(\nabla_{\mu}T^{\mu}_{~{}r})=0$ gives $V^{\prime}$ in terms of $(H,W,V,H^{\prime})$, thereby allowing one to eliminate $V^{\prime}$ from $W^{\prime}$ and $H^{\prime\prime}$.
@@
$W^{\prime}=\alpha_{W,H^{\prime}}H^{\prime}+\alpha_{W,H}H+\alpha_{W,W}W+\alpha_{W,V}V,$ (45)
@@ B.1.6 Electric-type tidal deformability
$\left.\frac{\mathbb{I}_{H}}{\mathbb{E}_{H}}=-\frac{RH_{T}^{\prime}-y_{\rm in}^{+}H_{T}}{RH_{R}^{\prime}-y_{\rm in}^{+}H_{R}}\right|_{r=R}.$ (57)
@@
$\bar{\lambda}_{\ell}=\frac{2}{\left(2\ell-1\right)!!}\frac{1}{{\cal C}^{2\ell+1}}\frac{p_{\ell,0}^{+}+p_{\ell,1}^{+}{\cal C}+p_{\ell,2}^{+}{\cal C}^{2}+p_{\ell,3}^{+}{\cal C}^{3}}{q_{\ell,0}^{+}+q_{\ell,1}^{+}{\cal C}+q_{\ell,2}^{+}{\cal C}^{2}+q_{\ell,3}^{+}{\cal C}^{3}},$ (58)
@@ B.2.1 Stellar perturbation equations
$h_{\mu\nu}^{-}dx^{\mu}dx^{\nu}=\left(2h^{\ell m}S_{\theta}^{\ell m}dtd\theta+2h^{\ell m}S_{\varphi}^{\ell m}dtd\varphi\right.$
$\left.+2h_{1}^{\ell m}S_{\theta}^{\ell m}drd\theta+2h_{1}^{\ell m}S_{\varphi}^{\ell m}drd\varphi\right)e^{-i\omega t}.$ (60)
@@
In the odd-parity sector, the metric perturbation is not coupled to the pressure and density perturbations but induces the displacement of the four velocity. The Eulerian perturbation of the four velocity is parametrized by a function $U=U(r)$ Price and Thorne (1969); Pani et al. (2015b, 2018)
$\delta u^{t}=$ $0,\quad\delta u^{r}=0,$ (61)
$\delta u^{A}=$ $\frac{Ue^{\nu/2}}{4\pi r^{2}\left(\rho+p\right)}\left(S_{\theta}^{\ell m},\frac{S_{\varphi}^{\ell m}}{\sin^{2}\theta}\right)e^{-i\omega t}.$ (62)
@@
One finds from $\delta(\nabla_{\mu}T^{\mu}_{~{}\theta})=0$ that